A body is projected at an angle of 30° to the horizontal with a velocity of 150ms-1 . calculate the greatest height (g =10)


1
Expert's answer
2021-11-17T10:40:03-0500

Given:

"v_0=150\:\rm m/s"

"\theta=30^{\circ}"

"g=10\:\rm m/s^2"


The maximum height

"h_{\max}=\frac{v_0^2\sin^2\theta}{2g}"

"h_{\max}=\frac{150^2\sin^2 30^{\circ}}{2*10}=281\:\rm m"